OpenStudy (anonymous):

Subtract, simplify if possible: v-8/v - 2v-13/5v

OpenStudy (saifoo.khan):

\[\Large \frac{v-8}{v}-\frac{2v-13}{5v}\] Like that?

OpenStudy (anonymous):

yep

OpenStudy (saifoo.khan):

take the LCM, \[\Large \frac{5v(v-8)-v(2v-13)}{5v(v)}\]

OpenStudy (anonymous):

you can multiply the first factor by 5, so that you then have a common denominator for both... Don't forget to also multiply the top line by 5. This gives you [5(v-8)/5v] - [2v-13 / 5v] and you can then use the same denominator, and subtract across the top line :)

OpenStudy (anonymous):

hmm feldy I think subtracting rational expressions is a little different than subtracting fractions. We have not had to do either of those things like you would with a regular fraction. I might be wrong though, it is all very confusing.

OpenStudy (anonymous):

Saifoo how did you find the LCM?

OpenStudy (anonymous):

OK my bad! Haven't dont this in a while haha

OpenStudy (anonymous):

its okay Im learning it for the first time. I am trying to use my book for help, however none of the example problems are quiet like this one.

OpenStudy (saifoo.khan):

do u know the basics of LCM? i

OpenStudy (anonymous):

we learned them last week, however I was nnot very solid on that part, so throwing me into this was a nightmare

OpenStudy (saifoo.khan):

Do u wanna learn?

OpenStudy (anonymous):

I would love too

OpenStudy (saifoo.khan):

Let's finsid your confusions! Come here!